OpenStudy (anonymous):

MY DUMB BUTT NEEDS HELP AGIN!!!!What is the solution to this equation? 2(x + 5) - 16 = 8 A. x = 12 B. x = 7 C. x = 2 D. x = 17

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

yes is b

OpenStudy (anonymous):

2(x + 5) - 16 = 8 + 16 +16 add 16 to both sides. 2(x +5) = 24 2x + 10 = 24 (distributive property) - 10 -10 subtract 10 from each side 2x = 14 /2 /2 divide each side by 2 x= 7...answer:)

OpenStudy (anonymous):

but is it be like a og

OpenStudy (anonymous):

yeah i just guessed then pluged in the numbers